ST IVES TOWN COUNCIL

PART-TIME SECRETARY/ CLERICAL ASSISTANT

Applications are invited for the above position in a small but busy office. The hours of work will be 1.00pm to 5.00pm, Monday to Thursday and 12.30pm to 4.30pm on Fridays (20 hours per week).

The successful candidate will support the Town Clerk and two other parttime office staff in carrying out the administrative functions of the Town Council.

Applicants should have a good working knowledge of Microsoft Office packages and be capable of undertaking a variety of office duties, including dealing with the general public.

The salary will commence at £11319 per annum pro rata.

Further details are available from: Anthony Watson, Town Clerk, St Ives Town Council, Town Hall, St Ives, Cambridgeshire PE17 4AN Tel: 01480 467505 To whom CVs should be sent by Monday 26th July 1999

CAMBRIDGE CITY COUNCIL

PERSONNEL DEPARTMENT

The Personnel Section, Cambridge City Council, 10 Downing Street, Cambridge CB2 3DS Telephone : (01223) 458100

ARCHIVIST REF: HM69

COUNTY COUNCIL RECORDS AND REGISTRY SERVICES (TEMPORARY CONTRACT, INITIALLY FOR THREE MONTHS, COMMENCING 2 AUGUST 1999) Based at: Heritage Services, Records Office, Shire Hall, Cambridge Grade: Scale 6 Salary: 16,719- 17,838 Hours: 37 per week We are looking for an archivist with some experience of records management

and registry systems to manage our records services to the County Council during the absence on secondment of our senior archivist. If you have a Diploma in Archive Administration or equivalent, some experience of current records work, an interest in developing ITC systems and are hard-working, with a handson approach, this could be the opportunity you are looking for to develop your skills. For further information contact Liz Whittingham, Archivist in Charge on 01223 717295. Apply by application form only, available with further details from The Human Resources Team, Education, Libraries & Heritage, Box No ELH1403, Castle Court, Shire Hall, Cambridge, CB3 OAP or telephone 01223 717800 (24 hour answerphone). Closing date 21 July 1999. Interviews will be held on 23 July 1999.